Divide and simplify. (3z-12)/(14) / (z-4)/(77z)
Found 2 solutions by chiexpert, stanbon:
Answer by chiexpert(48)

Divide and simplify. (3z-12)/(14) / (z-4)/(77z)
Here's the Rule for division...
To divide fractions, convert the division process to a multiplication process by using the following steps.
1. Change the "÷" sign to "x" and invert the fraction to the right of the sign.
2. Multiply the numerators.
3. Multiply the denominators.
4. Re-write your answer in its simplified or reduced form, if needed
Once you complete Step #1 for dividing fractions, the problem actually changes from division to multiplication.
example:
1/2 ÷ 1/3 = 1/2 x 3/1
1/2 x 3/1 = 3/2
Simplified Answer is 1 1/2
Now find your answer

Answer by stanbon(75887)

Divide and simplify. [(3z-12)/(14)] / [(z-4)/(77z)]
--------------------------
Invert the denominator and multiply:
---
= [(3(z-4)/(2*7)]
---
cancel any factor common to numerator and denominator
such as (z-4) and 7:
----
= [3/2]
----
= (33z/2]
=============
Cheers,
Stan H.